### v1

@@ Abstract
Urdu, as a low-resource language, lacks effective semantic content recommendation systems, particularly in the domain of personalized news retrieval. Existing approaches largely rely on lexical matching or language-agnostic techniques, which struggle to capture semantic intent and perform poorly under varying query lengths and information needs. This limitation results in reduced relevance and adaptability in Urdu content recommendation. We propose ULTRA (Urdu Language Transformer-based Recommendation Architecture), an adaptive semantic recommendation framework designed to address these challenges. ULTRA introduces a dual-embedding architecture with a query-length aware routing mechanism that dynamically distinguishes between short, intent-focused queries and longer, context-rich queries. Based on a threshold-driven decision process, user queries are routed to specialized semantic pipelines optimized for either title/headline-level or full-content/document level representations, ensuring appropriate semantic granularity during retrieval. The proposed system leverages transformer-based embeddings and optimized pooling strategies to move beyond surface-level keyword matching and enable context-aware similarity search. Extensive experiments conducted on a large-scale Urdu news corpus demonstrate that the proposed architecture consistently improves recommendation relevance across diverse query types. Results show gains in precision above 90% compared to single-pipeline baselines, highlighting the effectiveness of query-adaptive semantic alignment for low-resource languages. The findings establish ULTRA as a robust and generalizable content recommendation architecture, offering practical design insights for semantic retrieval systems in low-resource language settings.
@@
Urdu Content Recommendation System, Intelligent Information Retrieval, Semantic Search, Urdu NLP, Dual-Embedding Framework, Query Length Optimization,
@@ I Introduction
Text-based content recommendation systems have become indispensable in the digital era, transforming how users discover and consume information amidst an overwhelming volume of content [1, 2]. By leveraging user behavior, interests, and search queries, these systems curate personalized content feeds, enhancing engagement and information accessibility. At their core, modern systems rely on sophisticated semantic understanding to transcend simple keyword matching, aiming to grasp user intent and the contextual meaning of content for accurate alignment [3, 4]. Figure 2 illustrates the evolution of text-based recommendation and retrieval techniques from traditional lexical methods to modern semantic approaches.
@@
This challenge of semantic recommendation is not uniform; it is heavily modulated by linguistic and computational resource disparities [5]. For high-resource languages like English, advanced transformer-based architectures have achieved remarkable success [6, 7, 8]. However, for linguistically rich but computationally low-resource languages such as Urdu, the landscape is starkly different. Urdu, with large number of speakers worldwide and written in the complex Nastaliq script, presents unique morphological and syntactic challenges [9, 10]. Its context sensitivity, diglossia, and significant code-mixing, particularly with English and Roman Urdu in digital spaces, complicate automated semantic analysis [11, 12, 13]. This creates a critical barrier to accessing relevant digital content for a vast population.
@@
A principal drawback of these existing approaches is their rigid, one-size-fits-all architecture. They typically process all user queries through a single, uniform matching pipeline, whether against headlines, full articles, or entity graphs [19]. This design fails to account for the inherent diversity in user query patterns, which range from short, intent-focused searches (e.g., ”result of elections”) as in Figure 2,
Figure 2: Example of a short, intent-focused Urdu query
to long, descriptive inquiries seeking specific contextual details (e.g.,“regarding the results of Pakistan’s 2024 national elections, which political parties secured how many seats in different provinces and constituencies?”) as shown in Figure 3. [20].
Figure 3: Example of a long, context-rich Urdu query
@@
Consequently, a concise query may be inadequately matched against verbose article bodies, losing precision, while a detailed query may not fully leverage the rich semantics embedded in full-text content, compromising recall [21]. Furthermore, scalability remains a pervasive concern, as the computational cost of performing real-time similarity searches on high-dimensional transformer embeddings across large and growing content corpora is often prohibitive, limiting practical deployment [22, 23]. There is a clear, unaddressed need for an adaptive system that dynamically tailors its retrieval strategy to the nature of the user’s query [27, 28, 29].
@@
To address these critical limitations, we propose ULTRA (Urdu Language Transformer Based Recommendation Architecture), a novel dual-pathway framework designed for adaptive, high-precision semantic text-based content recommendation in Urdu. The core innovation of ULTRA is an intelligent routing mechanism that dynamically directs user queries to specialized semantic pipelines based on query characteristics. This ensures the granularity of the user’s informational need is aligned with the most appropriate level of document representation. By leveraging state-of-the-art transformer models fine-tuned for Urdu within an optimized vector search infrastructure, ULTRA achieves genuine semantic comprehension beyond lexical coincidence, while maintaining efficiency for large-scale deployment [29].
II Our Novel Contributions Explained through Use Cases
@@
Figure 4 provides an abstract representation of the ULTRA architecture, depicting the structural organization and functional relationships between its core components. Building upon this architectural foundation, the main contributions of this work can be summarized as follows:
II-A Use Cases in Text Based Content Recommendation Systems
@@
When recommending text-based content on content-based recommendation systems such as news websites, e-commerce sites, blog platforms, and digital libraries, two primary use cases emerge: (1) users enter a query into a search bar, triggering recommendations aligned with the query; and (2) users are engaged with a specific webpage, such as reading a news article, blog post, or product description on an e-commerce site, where related content is suggested below the current page.
@@
Our proposed architecture accommodates both scenarios. In the case of a user entering a query via the search bar—typically consisting of keywords and thus classified as a short query—the input is routed to a pipeline optimized for short queries. Conversely, when recommending content based on the webpage the user is currently viewing, the text on that specific web page serves as the query, which falls under the long query category. Recommendations are then generated relative to the topic the user is actively engaged with. However users can also type long queries in search bar; such inputs would similarly be directed to the long-query-optimized pipeline. To distinguish between short and long queries, we establish a predefined threshold value. The methodology for selecting and determining this threshold is detailed in section Section IV-A2.
II-B Dual Pipeline Approach
@@
The distinction between short and long queries is motivated by experimental findings detailed in Section V. Our evaluation revealed that short queries, composed of a few keywords, exhibit poor precision when matched against embeddings of lengthy content (e.g., complete articles or blog posts). However, matching these queries against headline embeddings, which are comparable in length and linguistic structure, yielded substantially higher precision.
@@
These observations drove our dual-pipeline architectural design. Rather than matching compact queries directly against full document embeddings, we propose an alternative strategy: (1) encode content semantics via headline embeddings, (2) retrieve relevant headlines based on query-headline similarity, and (3) recommend the complete documents associated with top-ranked headlines. This approach exploits the semantic and structural correspondence between query brevity and headline conciseness, thereby improving recommendation accuracy for Urdu text corpora.
II-C Major Points
@@
  • We propose ULTRA, a novel adaptive recommendation architecture that introduces a query-length threshold-based routing mechanism, as established in Section II-A and II-B, to dynamically optimize semantic matching, a novel approach for Urdu text-based content recommendation.
@@
  • We conduct an extensive empirical evaluation of embedding generation and pooling strategies, establishing optimal configurations for representing both title-level and full document-level Urdu content.
@@
  • We implement and rigorously compare multiple dimensionality reduction techniques as well as reduced dimension sizes, identifying the most effective method and dimension size for compressing semantic embeddings to enable efficient, large-scale similarity search without significant fidelity loss.
@@
  • We demonstrate through comprehensive experiments that ULTRA sets a new performance benchmark, consistently achieving precision rates exceeding 90% and significantly outperforming existing non-adaptive semantic and lexical baselines for Urdu text based content recommender systems.
Figure 4: Architecture diagram of proposed ULTRA recommendation framework
@@ III Related Work
Our ULTRA framework addresses these gaps by introducing a dual-pathway system with query routing, optimized pooling, dimensionality reduction, and vector databases. This advances Urdu text-based content recommendation, generalizing beyond news to handle diverse textual corpora efficiently. Table I outlines the key limitations of prior work in Urdu text-based content recommendation systems and illustrates how our proposed ULTRA framework addresses them.
TABLE I: Comparative analysis of urdu text-based content recommendation systems and the proposed ULTRA framework
Ref. Year Model/Technique Dataset Paradigm Key Contributions Identified Limitations
Kanwal et al. [14] 2021 Literature Review / Survey N/A (Review Paper) Survey of Text-based RS Comprehensive review of methods; Identified core challenges for low-resource languages like Urdu. No implementation; highlights gaps but provides no solution.
Abbas et al. [16] 2022 TF-IDF + Cosine Similarity, BERT Embeddings Custom Urdu News (1,160 articles) Content-based Filtering Introduced semantic BERT embeddings for Urdu news, moving beyond pure lexical matching. Limited dataset; single-pipeline design; no adaptation for different query types; relies partly on lexical matching.
Kanwal et al. [17] 2022 Entity Linking + RNN Hybrid with TransE SKRS Dataset (16,738 articles) Knowledge Graph-based Hybrid Pioneered Urdu Wikification; used entity knowledge graph to enrich semantics for recommendation. Moderate accuracy (60.8%); complex pipeline; performance tied to entity presence; no query adaptation.
Kanwal et al. [24] 2024 Entity-Enriched Framework Expanded Dataset (23,250 articles) Entity-Based Semantic Recommendation Advanced entity-based matching, exploiting hidden semantic features for improved precision. Heavy reliance on entity recognition; no query-adaptive routing; limited scalability for real-time systems.
@@ IV Proposed Methodology
In this research, we propose a methodology to develop a robust Urdu news recommendation system using semantic search powered by transformer-based embeddings. Figure 4 illustrates the architecture of the proposed system, and its modules are described below.
@@ IV-A Problem Formulation
Given a user query in Urdu and a corpus of Urdu news articles, the objective is to retrieve the $k$ most semantically relevant articles. We formalize this as a semantic search problem in a high-dimensional vector space.
Let the article corpus be a set $D=\{d_{1},d_{2},\ldots,d_{m}\}$, where each article $d_{i}\in D$ is a tuple comprising its title and full textual content: $d_{i}=(h_{i},c_{i})$, with $h_{i}$ headline and $c_{i}$ the content/complete news artcile corresponding to that headline.
A user query is denoted by $q\in\mathcal{Q}$, where $\mathcal{Q}$ is the space of all possible Urdu text queries. The character length of a query is given by the function $\ell(q)$.
The core task of the system is to retrieve a set of articles that are semantically relevant to a user’s query. We model this using a binary relevance function $\text{rel}:\mathcal{Q}\times D\rightarrow\{0,1\}$. For a given query $q$ and article $d_{i}$, a human evaluator assigns:
$\text{rel}(q,d_{i})=\begin{cases}1,&\text{if article $d_{i}$ is relevant to query $q$,}\\ 0,&\text{otherwise.}\end{cases}$
A relevant article returned for a query constitutes a True Positive (TP).
Operationally, the system employs cosine similarity in a learned embedding space as an efficient proxy to approximate semantic relevance. It retrieves the top-$k$ articles, denoted $R_{q}^{k}$, ranked by this similarity score:
$R_{q}^{k}=\underset{d_{i}\in D,|R|=k}{\arg\max}\;\cos(\mathbf{v}^{\prime}_{q},\mathbf{v}^{\prime}_{i}).$ (1)
The primary performance metric is precision. For a single query $q$, the precision at $k$ (Precision@k) is the proportion of retrieved articles that are relevant:
@@ IV-A2 Query-adaptive routing
$\text{Retrieval Space}=\begin{cases}\{\mathbf{v}_{h_{i}}\}_{i=1}^{m},&\ell(q)<\theta\\ \{\mathbf{v}_{c_{i}}\}_{i=1}^{m},&\ell(q)\geq\theta\end{cases}\ \ \begin{array}[]{l}\text{(Short Query Pathway)}\\ \text{(Long Query Pathway)}\end{array}$
@@ IV-A4 Similarity computation and ranking
This cosine score serves as the operational retrieval metric. The final ranked list $R_{q}^{k}$ is obtained by sorting all articles in descending order by $\text{sim}(q,d_{i})$ and selecting the top $k$. It is important to note that while this geometric measure is an effective proxy for semantic relatedness, the final assessment of recommendation quality is based on human-evaluated relevance, as defined in Section IV-A. The system’s objective is to maximize the mean precision by optimizing the underlying embedding generation, routing, and dimensionality reduction components to produce $R_{q}^{k}$ that aligns as closely as possible with human relevance judgments.
This formulation provides a rigorous mathematical foundation for the subsequent development and evaluation of the ULTRA framework’s components, including the selection of $P$ (pooling strategies), the determination of $\theta$ (routing threshold), the choice of $R$ (dimensionality reduction technique), and the optimization of $d^{\prime}$ (reduced dimension). Notably, this problem formulation is general and can be applied to any recommendation system, although we specify news articles in our implementation due to the availability of the largest Urdu corpus dataset for evaluation. Nonetheless, the framework extends to a wide array of recommendation systems, encompassing digital libraries, news recommenders, product recommenders, and blog page recommenders.

### v2

@@ Abstract
Urdu being a low resource language lacks effective semantic content recommendation systems especially in the area of personalized news retrieval. The current methods are mostly based on lexical matching or language-agnostic methods, which are not very effective at capturing semantic intent and do not work well when the query length and information requirements vary. This weakness leads to poor relevance and flexibility in Urdu content recommendation. To overcome these obstacles, we propose ULTRA (Urdu Language Transformer-based Recommendation Architecture) which is an adaptive semantic recommendation architecture. ULTRA presents a dual-embedding architecture, which has a query-length sensitive routing scheme that dynamically differentiates between short, intent-oriented queries and longer, context-rich queries. According to a threshold-based decision process, user queries are directed to either title/headline-level or full-content/document level-specific semantic pipelines, which guarantee the right level of semantic granularity during retrieval. The suggested system exploits the transformer-based embeddings and optimized pooling techniques to go beyond the surface-level matching of keywords and allow similarity search based on context. The large-scale experiments performed on a large Urdu news corpus indicate that the suggested architecture is always effective in enhancing the relevance of recommendations in relation to different query types. The findings demonstrate precision of over 90% as compared to single-pipeline baselines. This underscores the usefulness of query-adaptive semantic alignment to low-resource languages. The results make ULTRA a strong and scalable content recommendation structure, providing useful design knowledge to semantic retrieval systems on low-resource language environments.
@@
Content Recommenders, Urdu Content Recommendation System, Intelligent Information Retrieval, Semantic Search, Urdu NLP, Dual-Embedding Framework
@@ I Introduction
Text-based content recommendation systems have become indispensable in the digital era, transforming how users discover and consume information amidst an overwhelming volume of content [1, 2]. By leveraging user behavior, interests, and search queries, these systems curate personalized content feeds, enhancing engagement and information accessibility. At their core, modern systems rely on sophisticated semantic understanding to go beyond simple keyword matching, aiming to grasp user intent and the contextual meaning of content for accurate alignment [3, 4]. The evolution of text-based recommendation and retrieval techniques from traditional lexical methods to modern semantic approaches is illustrated in Figure 1.
@@
This challenge of semantic recommendation is not uniform; it is heavily modulated by linguistic and computational resource disparities [5]. For high-resource languages like English, advanced transformer-based architectures have achieved remarkable success [6, 7, 8]. However, for linguistically rich but computationally low-resource languages such as Urdu, the landscape is clearly different. Urdu, with large number of speakers worldwide and written in the complex Nastaliq script, presents unique morphological and syntactic challenges [9, 10]. Its context sensitivity, diglossia, and significant code-mixing, particularly with English and Roman Urdu in digital spaces, complicate automated semantic analysis [11, 12, 13]. This creates a critical barrier to accessing relevant digital content for a vast population.
@@
A principal drawback of these existing approaches is their rigid, one-size-fits-all architecture. They typically process all user queries through a single, uniform matching pipeline, whether against headlines, full articles, or entity graphs [19]. This design fails to account for the inherent diversity in user query patterns, which range from short, intent-focused searches as in Figure 2,
Figure 2: Example of a short, intent-focused Urdu queries
to long, descriptive inquiries seeking specific contextual details as shown in Figure 3. [20].
Figure 3: Example of a long, context-rich Urdu queries
@@
Consequently, a concise query may be inadequately matched against verbose article bodies, losing precision, while a detailed query may not fully leverage the rich semantics embedded in full-text content, compromising recall [21]. Furthermore, scalability remains a pervasive concern. The computational cost of performing real-time similarity searches on high-dimensional transformer embeddings across large and growing content corpora limits practical deployment [22, 23]. There is a clear need for an adaptive system that dynamically tailors its retrieval strategy to the nature of the user’s query [27, 28, 29].
@@
In order to overcome these severe constraints, we introduce ULTRA (Urdu Language Transformer Based Recommendation Architecture), a new dual-pathway architecture that is used to recommend semantic text-based content in Urdu with a high level of accuracy and adaptability. The main innovation of ULTRA is a smart routing system that dynamically routes user queries to specialized semantic pipelines depending on query properties. This makes sure that the granularity of the informational need of the user is matched with the most suitable level of document representation. With the help of state-of-the-art transformer models optimized to Urdu in an optimized vector search infrastructure, ULTRA can achieve true semantic understanding beyond lexical coincidence, and is efficient enough for large scale use [29].
II Key Contributions
@@
An abstract representation of the ULTRA architecture, depicting the structural organization and functional relationships between its core components is presented in Figure 4. Building upon this architectural foundation, the main contributions of this work can be summarized as follows:
II-A Use Cases in Text Based Content Recommendation
@@
When recommending text-based content on content recommendation systems such as news websites, e-commerce sites, blog platforms, and digital libraries, two primary use cases emerge: (1) users enter a query into a search bar, triggering recommendations aligned with the query; and (2) users are engaged with a specific webpage, such as reading a news article, blog post, or product description on an e-commerce site, where related content is suggested below the current page.
@@
Our proposed architecture accommodates both scenarios. In the case of a user entering a query via the search bar that typically consists of keywords and thus classified as a short query, the input is routed to a pipeline optimized for short queries. Conversely, when recommending content, based on the webpage the user is currently viewing, the text on that specific web page serves as the query, which falls under the long query category. Recommendations are then generated relative to the topic the user is actively engaged with. However users can also type long queries in search bar; such inputs would similarly be directed to the long-query-optimized pipeline. To distinguish between short and long queries, we establish a predefined threshold value. The methodology for selecting and determining this threshold is detailed in section Section IV-A2.
II-B Proposed Dual-Pipeline Solution for Identified Use Cases
@@
The use cases described in Section II-A motivate the need for different processing strategies for short and long queries. Our experimental analysis in Section V revealed that short queries, composed of a few keywords, exhibit poor precision when matched against embeddings of lengthy content (e.g., complete articles or blog posts). However, matching these queries against headline embeddings, which are comparable in length and linguistic structure, yielded substantially higher precision.
@@
These observations drove our dual-pipeline architectural design. Rather than matching compact queries directly against full document embeddings, we propose an alternative strategy: (1) encode content semantics via headline embeddings, (2) retrieve relevant headlines based on query-headline similarity, and (3) recommend the complete documents associated with top-ranked headlines. This approach exploits the semantic and structural correspondence between query brevity and headline conciseness. Hence, it improves recommendation precision for Urdu text corpora.
II-C Summary of Main Contributions
@@
  • We developed ULTRA, a novel adaptive recommendation architecture that introduces a query-length threshold-based routing mechanism, as established in Section II-A and II-B, to dynamically optimize semantic matching, a novel approach for Urdu text-based content recommendation.
@@
  • We conduct empirical evaluation of embedding generation and pooling strategies, establishing optimal configurations for representing both title-level and full document-level Urdu content.
@@
  • We implement and rigorously compare multiple dimensionality reduction techniques as well as reduced dimension sizes. This helps us identify the most effective method and dimension size for compressing semantic embeddings to enable efficient as well as large-scale similarity search without significant fidelity loss.
@@
  • We demonstrate through experiments that ULTRA sets a new performance benchmark, consistently achieving precision rates exceeding 90%. It also significantly outperforms existing non-adaptive semantic and lexical baselines for Urdu text based content recommender systems.
Figure 4: Architecture diagram of proposed ULTRA framework
@@ III Related Work
Our ULTRA framework addresses these gaps by introducing a dual-pathway system with query routing, optimized pooling, dimensionality reduction, and vector databases. This advances Urdu text-based content recommendation, generalizing beyond news to handle diverse textual corpora efficiently. Table I outlines the key limitations of prior work in Urdu text-based language processing systems and illustrates how our proposed ULTRA framework addresses them.
TABLE I: Comparative analysis of foundational Urdu NLP systems and their contributions to text-based content recommendation
Ref. Yr. Model/Technique Dataset Paradigm Key Contributions Identified Limitations
[14] 2021 Literature Review / Survey N/A (Review Paper) Survey of Text-based RS Comprehensive review of methods; Identified core challenges for low-resource languages like Urdu. No implementation; highlights gaps but provides no solution.
[16] 2022 TF-IDF + Cosine Similarity, BERT Embeddings Custom Urdu News (1,160 articles) Content-based Filtering Introduced semantic BERT embeddings for Urdu news, moving beyond pure lexical matching. Limited dataset; single-pipeline design; no adaptation for different query types; relies partly on lexical matching.
[17] 2022 Entity Linking + RNN Hybrid with TransE SKRS Dataset (16,738 articles) Knowledge Graph-based Hybrid Pioneered Urdu Wikification; used entity knowledge graph to enrich semantics for recommendation. Moderate accuracy (60.8%); complex pipeline; performance tied to entity presence; no query adaptation.
[24] 2024 Entity-Enriched Framework (SEEUNRS) Expanded Dataset (23,250 articles) Entity-Based Semantic Recommendation Advanced entity-based matching, exploiting hidden semantic features for improved precision. Heavy reliance on entity recognition; no query-adaptive routing; limited scalability for real-time systems.
[50] 2023 Deep Learning for Urdu News Classification Urdu News 1M Dataset (1M articles) Text Classification Large-scale Urdu news classification using transformers; demonstrated scalability. Classification-focused; no personalization or recommendation components; single-representation approach.
[51] 2021 RoBERTa-based Pre-trained Language Model Urdu news corpus Language Model Pre-training Earliest public RoBERTa-based model for Urdu; handles right-to-left Nastaliq script, morphology, and semantics; foundational for Urdu semantic NLP tasks. Focused on language modeling only; no recommendation-specific fine-tuning or retrieval optimization.
[52] 2023 Multilingual Sentence-BERT for Indic Languages Synthetic NLI/STS Cross-lingual Embeddings Extended sentence transformers to low-resource Indic languages; improved semantic similarity. General-purpose embeddings; not optimized for recommendation tasks; no query routing mechanism.
[53] 2026 U-RR²: Two-stage framework integrating TF-IDF, dense embeddings, and SVMrank re-ranking with feature ensemble CURE, ROSHNI, UIR_21 benchmarks Two-Stage Dense Retrieval & Re-ranking Proposes a hybrid two-stage Urdu document retrieval framework with large-scale word embeddings, achieving state-of-the-art performance on Urdu IR benchmarks (MAP 0.78–0.81, P@10 0.79–0.82, $F_{1}$ up to 0.90). Designed for ad-hoc document retrieval rather than personalized semantic recommendation; lacks query-adaptive routing, granularity differentiation, recommendation-specific optimizations (pooling strategies, dimensionality reduction), and scalable real-time personalization mechanisms.
@@ IV Proposed Methodology
In this research, we propose a methodology to develop a robust Urdu news recommendation system using semantic search powered by transformer-based embeddings. The architecture of the proposed system is illustrated in Figure 4, and its modules are described below.
@@ IV-A Problem Formulation
Considering a user query in Urdu and a collection of Urdu news stories, the task is to find the top $k$ semantically relevant news stories. We state this as a semantic search problem in high-dimensional vector space.
Suppose the article corpus is a set $D=\{d_{1},d_{2},\ldots,d_{m}\}$ where each article $d_{i}$ is a pair: $d_{i}=(h_{i},c_{i})$ where $h_{i}$ is the headline and $c_{i}$ is the content/complete news article that is the subject of that headline.
The query made by a user is represented by $q\in\mathcal{Q}$ with the space of all possible Urdu text queries denoted by $\mathcal{Q}$. The length of a query is provided by the function $\ell(q)$.
The main activity of the system is to access a collection of articles that are semantically related to a query of a user. We represent this by a binary relevance function, which we denote as $\text{rel}:\mathcal{Q}\times D\rightarrow\{0,1\}$. Given a query $q$ and an article $d_{i}$, a human evaluator assigns:
$\text{rel}(q,d_{i})=\begin{cases}1,&\text{if article }d_{i}\text{ is relevant to query }q,\\ 0,&\text{otherwise.}\end{cases}$
A True Positive (TP) is a relevant article that has been returned in response to a query.
The system uses cosine similarity in learned embedding space as an effective proxy to estimate semantic relevance operationally. It returns the first $k$ articles, denoted $R_{q}^{k}$, in terms of this similarity score:
$R_{q}^{k}=\underset{d_{i}\in D,|R|=k}{\operatorname{argmax}}\cos(\mathbf{v}_{q},\mathbf{v}_{i}).$ (1)
Precision is the key performance indicator. For a single query $q$, the precision at $k$ (Precision@k) is the proportion of retrieved articles that are relevant:
@@ IV-A2 Query-adaptive routing
$\text{Retrieval Space}=\begin{cases}\{\mathbf{v}_{h_{i}}\}_{i=1}^{m},&\ell(q)<\theta\\ \{\mathbf{v}_{c_{i}}\}_{i=1}^{m},&\ell(q)\geq\theta\end{cases}\ \ {\small\begin{array}[]{l}\text{(Short Query Route)}\\ \text{(Long Query Route)}\end{array}}$
@@ IV-A4 Similarity computation and ranking
This cosine score serves as the operational retrieval metric. The final ranked list $R_{q}^{k}$ is obtained by sorting all articles in descending order by $\text{sim}(q,d_{i})$ and selecting the top $k$. It is important to note that while this geometric measure is an effective proxy for semantic relatedness, the final assessment of recommendation quality is based on human-evaluated relevance, as defined in Section IV-A. This system aims to maximize the average precision by optimizing the underlying embedding generation, routing, and dimensionality reduction modules to generate $R_{q}^{k}$ that is as close as possible to human relevance judgments.
This formulation gives a strict mathematical basis to the further development and analysis of the components of the ULTRA framework, such as the selection of $P$ (pooling strategies), the determination of $\theta$ (routing threshold), the selection of $R$ (dimensionality reduction technique) and optimization of $d$ (reduced dimension). It is important to note that this problem formulation is universal and can be used with any recommendation system, but in our implementation, we use news articles because of the availability of the largest dataset of Urdu corpus to evaluate. However, the framework is applicable to a broad range of recommendation systems, which include digital libraries, news recommenders, product recommenders, and blog page recommenders.